OpenStudy (anonymous):

vectors a=<3,2> b=<2,-1>, c=<7,1> Find value of s and t, where c=sa+tb. How do I start this?

OpenStudy (anonymous):

sa=<3s,2s>, and tb=<2t,-1t> ??

OpenStudy (anonymous):

so you now have a system of equations: 3s + 2t = 7 and 2s - 1t = 1

OpenStudy (anonymous):

ill try that.. but set 3s+2t=|c| guess its not right

OpenStudy (anonymous):

ye its correct =)

OpenStudy (anonymous):

thanks

OpenStudy (anonymous):

no problem

OpenStudy (anonymous):

but im still couldnt see how you can set them equal to something. like 3s+2t=7

OpenStudy (anonymous):

is there rule or? cuz i use <i,j,k> so when adding, its all is, js, ks. like <3s+2t,2s-t>=|c|, but it turned out, cant be solved

OpenStudy (anonymous):

they are the coefficients of the components 3<xs, ys> + 2 <xt + yt> = <xc, yc>

OpenStudy (anonymous):

so 3 xs + 2 xt = xc and 3ys ++ 2yt = yt ** xs being the x component of s and ys being the y component etc

OpenStudy (anonymous):

sorry that second one is meant to read 3ys + 2yt = yc

OpenStudy (anonymous):

k, make sense. Thanks again